MEGHÍVÓ

2002. november 4-én, hétfőn, a Nagyteremben 14 órakor kezdődő Összintézeti Szemináriumunk programja:

MARK HOGARTH (Cambridge, UK)

A NEW CONCEPT OF COMPUTABILITY

című előadása.

This talk is about so-called SAD_{n} computers, a class of Turing machinebased computers that utilize the geometry of Malament-Hogarth spacetimes to perform non-Turing computable feats. It will be shown how, for example, SAD_{n} can decide n-quantifier arithmetic but not (n + 1)-quantifier arithmetic, a result which reveals how neatly the SAD_{n} family maps into the Kleene arithmetical hierarchy. I will argue that these investigations give rise to a radically new concept of computability, one that accounts for both Turing’s machine and these new non-Turing machines. Intriguingly this new concept makes no appeal to any kind of Church-Turing conjecture.

MINDEN ÉRDEKLŐDŐT SZERETETTEL VÁRUNK!

2002. november 4-én, hétfőn, a Nagyteremben 14 órakor kezdődő Összintézeti Szemináriumunk programja:

MARK HOGARTH (Cambridge, UK)

A NEW CONCEPT OF COMPUTABILITY

című előadása.

This talk is about so-called SAD_{n} computers, a class of Turing machinebased computers that utilize the geometry of Malament-Hogarth spacetimes to perform non-Turing computable feats. It will be shown how, for example, SAD_{n} can decide n-quantifier arithmetic but not (n + 1)-quantifier arithmetic, a result which reveals how neatly the SAD_{n} family maps into the Kleene arithmetical hierarchy. I will argue that these investigations give rise to a radically new concept of computability, one that accounts for both Turing’s machine and these new non-Turing machines. Intriguingly this new concept makes no appeal to any kind of Church-Turing conjecture.

MINDEN ÉRDEKLŐDŐT SZERETETTEL VÁRUNK!